Can I please get full working and explanations for each step for this past exam question. Both parts (a) and (b) and (c). Thanks will up-voteQuestion B1 The isomerisation shown below has an equilibrium constant of K = 0.74 k1 trans-Co(en)2(H2O)OH) k1 cis-Co(en)2(H2O
Answer #1

(a)

rofun,(mpont K: pay- t-o 20-74 a tet a-x da dt t(a-t)-x ect e 0 dt &,(a-x.) dx lan)-A(a-) k,a (x) dt de e dk t a dt e ln Me t

(b)

kobs. = k1 + k-1 = 5.2×10-3 s-1

Half life of first order reaction ,t1/2 =( ln2) /kobs. = 0.693/(5.2×10-3) s = 133.298 s

= 133.3 s

(c)

(k_1 + k1 )×t = ln(xe/(xe-x))

For nearly completion reaction, k1 >>>k-1

k1 + k-1 = k1   and xe = [cis]0 and xe -x = [cis]

k1 t = ln([cis]0/[cis])

The concentration of cis-isomer decreases exponentially.

[cis]t = [cis]0e-k1t